OpenStudy (anonymous):

Can someone explain to me how to do the substitution method using this problem? Solve the given system, using the substitution method. y = 3x – 7 6x – 2y = 12 A. There is no solution. B. (12, 14) C. There are an infinite number of solutions. D. (14, 12)

OpenStudy (godlovesme):

substitute "3x-7" for y in the eqn "6x-2y=12" make sense?

OpenStudy (godlovesme):

@simsim789

OpenStudy (anonymous):

i don't really get it

OpenStudy (anonymous):

how do you "substitute" it? @Godlovesme

OpenStudy (godlovesme):

k y=3x-7 so 6x-2y=12 6x-2(3x-7)=12 can u solve for x?

OpenStudy (godlovesme):

@simsim789

OpenStudy (anonymous):

yah, so first you have to use the distribution property so 2*3x= 6x and 2*7=14 so now it would be 6x - 6x - 14=12 then you have to subtract the first 6x from the equation, am i doing it right? @Godlovesme

OpenStudy (godlovesme):

yes!! :D

OpenStudy (anonymous):

so now it would be -6x - 14 = -6x + 12 right?

OpenStudy (godlovesme):

wasn't it like this 6x-6x-14= 12 ???

OpenStudy (anonymous):

yah, but after you subtract the 6x you would get -6x - 14= 6x +12? i rlly suck at this stuff

OpenStudy (godlovesme):

aww it's alright :)) u don't need to subtract 6x cuz we don't have like erms on the other side  does that make sense?